OpenStudy (anonymous):

What is the lim as h ->0 when f(x)=

OpenStudy (anonymous):

\[f(x)= \left(\begin{matrix}\sqrt{(3+h)-\sqrt{3}} \\ h\end{matrix}\right)\]

OpenStudy (anonymous):

that's divided over h i dont know why it isnt showing.

OpenStudy (ash2326):

\[\lim_{h\to 0} \frac{\sqrt {(3+h)-\sqrt 3}}{h}\]

OpenStudy (anonymous):

that's what I intended but the root is over 3+h then - root of 3 all over h

OpenStudy (ash2326):

\[\lim_{h\to 0} \frac{\sqrt {(3+h)}-\sqrt 3}{h}\] You mean this?

OpenStudy (anonymous):

Yes. Thanks. That looks better.

OpenStudy (anonymous):

I need to rationalize it, right?

OpenStudy (ash2326):

Yes :) Multiply and divide by \(\sqrt {3+h}+ \sqrt 3\)

OpenStudy (anonymous):

I get to the point of (3+h-9)/(h(root(3+h)+root(3))) and then I messed up somewhere afterwards

OpenStudy (ash2326):

Ok, I'll illustrate \[\lim_{h\to 0} \frac{\sqrt {(3+h)}-\sqrt 3}{h}\times \frac{\sqrt {3+h}-\sqrt 3}{h}\] \[\lim_{h \to 0}\frac{(\sqrt {3+h})^2-(\sqrt 3)^2}{h\times (\sqrt{3+h}+\sqrt 3)}\] \[\lim_{h\to 0} \frac {3+h-3}{h\times (\sqrt {3+h}+\sqrt 3)}\] Can you take it from here?

OpenStudy (ash2326):

Sorry I messed the first step

OpenStudy (ash2326):

It's actually this \[\lim_{h\to 0} \frac{\sqrt {(3+h)}-\sqrt 3}{h}\times \frac{\sqrt {3+h}+\sqrt 3}{\sqrt {3+h}+\sqrt 3}\]

OpenStudy (anonymous):

Well, Would the limit then be 0? And I see that I went and multiplied the 3's in the first step LOL. fail.

OpenStudy (ash2326):

put h=0, here after simplifying \[\lim_{h\to 0} \frac {3+h-3}{h\times (\sqrt {3+h}+\sqrt 3)}\]

OpenStudy (anonymous):

Wait, I don't know how to simplify that.

OpenStudy (ash2326):

\[\lim_{h\to 0} \frac {3+h-3}{h\times (\sqrt {3+h}+\sqrt 3)}\] \[\lim_{h\to 0} \frac {h}{h \times(\sqrt {3+h}+\sqrt 3}\] \[\lim_{h\to 0} \frac {h}{h \times(\sqrt {3+h}+\sqrt 3}\] \[\lim_{h\to 0} \frac {1}{\sqrt {3+h}+\sqrt 3}\] now put h=0 here

OpenStudy (anonymous):

Oh..but what happens when you add roots? (algebra skills=none :/ )

OpenStudy (ash2326):

Works just like normal no. s \[3+3=2\times 3\] so \[\sqrt 3+\sqrt 3=2\times \sqrt 3 \ or\ 2\sqrt3\]